# Adapting AI to the Moment: Understanding the Dynamics of Parent–AI Collaboration Modes in Real-Time Conversations with Children

### 29 Mar 2026 v1

@@ 7.4. Design Implications
Grounded in our findings, we derive design implications for flexible and effective parent–AI collaboration.
@@ 7.4.3. Rethinking AI Literacy in Parenting
Collaboration flexibility varied with parents’ AI literacy: in our study, high-literacy parents dynamically reconfigured functions, whereas low-literacy parents tended to rely on fixed collaboration modes. This suggests that parent–AI collaboration systems may impose excessive complexity and cognitive burden on low-literacy parents, highlighting the need for literacy-aware designs that adapt system complexity to parents’ capabilities.
However, in parenting contexts, AI literacy extends beyond technical proficiency (Long and Magerko, 2020; Ng et al., 2021) to include relational sensitivity—knowing when and how to involve AI without disrupting the parent–child relationship.
At the same time, overly simplified designs may constrain parents’ agency and hinder long-term learning. This underscores the importance of building on prior work on fostering AI literacy (Lee et al., 2021; Ng et al., 2024), and balancing immediate usability with the gradual development of parents’ AI literacy.
@@ 7.5. Limitations and Future Work
Finally, COMPASS focused on six common scaffolding functions identified in our formative study, which do not cover the full design space. While this scope is sufficient for our purpose of examining parents’ shifting collaboration needs, future work could extend parent–AI collaboration to provide richer, more personalized support.
@@ 8. Conclusion
We developed COMPASS, a research probe designed to provide flexible scaffolding for parents during parent-child conversations. Using COMPASS as a probe in a study with 21 parent-child pairs, we find that parent–AI collaboration evolves from backstage support to partner-style co-performance, shifting with contextual factors to balance cognitive and emotional regulation. We identify three parental strategies, namely parent-oriented, child-oriented, and relationship-oriented strategies, extending Parental Goal Theory and informing the design of future parental support systems that enable flexible and effective partnerships.

### 01 Apr 2026 v2

@@ 7.3. Design Implications
Grounded in our findings, we derive design implications for designing parent–AI systems that support flexible collaboration in parent-child interactions.
@@ 7.3.3. Rethinking AI Literacy in Parenting
Collaboration flexibility varied with parents’ AI literacy: high-literacy parents dynamically reconfigured functions, whereas low-literacy parents relied on fixed collaboration modes, suggesting that current systems may impose excessive complexity and cognitive burden on low-literacy parents. This highlights the need for literacy-aware designs that adapt system complexity to parents’ capabilities. In parenting contexts, however, AI literacy extends beyond technical proficiency (Long and Magerko, 2020; Ng et al., 2021) to include relational sensitivity—knowing when and how to involve AI without disrupting the parent–child relationship. At the same time, overly simplified designs risk constraining parents’ agency and hindering long-term learning. Building on prior work on fostering AI literacy (Lee et al., 2021; Ng et al., 2024), these findings point to balancing immediate usability with the gradual development of parents’ AI literacy.
7.3.4. Context-Aware Restraint in Sensitive Parenting Moments
Parents adapted AI use to context: in calm, cognitively oriented moments, they preferred structured support, whereas in emotionally intense situations, they avoided directive supports and relied on alignment and mediation. While prior HCI work highlights context-aware support (Ogata and Yano, 2004; Chen et al., 2023, 2025b), we further identify the need for context-aware restraint. AI should not only know when to step in but also hold back in “no-go” moments (e.g., avoiding rigid responses when children are sad). This aligns with “calm technology” (Weiser and Brown, 1996, 1997) and reframes deliberate AI silence as a form of positive intelligence that protects sensitive parent–child interactions.
@@ 7.4. Limitations and Future Work
Finally, COMPASS focused on six common AI support functions identified in our formative study, which do not cover the full design space. While this scope is sufficient for our purpose of examining parents’ shifting collaboration needs, future work could extend parent–AI collaboration to provide richer, more personalized support.
@@ 8. Conclusion
We developed COMPASS, a research probe designed to provide flexible AI support for parents during parent-child conversations. Using COMPASS as a probe in a study with 21 parent-child pairs, we find that parent–AI collaboration evolves from backstage support to partner-style co-performance, shifting with contextual factors to balance cognitive and emotional regulation. We identify three parental strategies, namely parent-oriented, child-oriented, and relationship-oriented strategies, extending Parental Goal Theory and informing the design of future parental support systems that enable flexible and effective partnerships.